OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

The revenue of a company is represented by r(x) and the expenses of the company are represented by e(x). What would you do to find the profit, p(x)?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

A. r(x) + e(x) = p(x) B. r(x) – e(x) = p(x) C. e(x) – r(x) = p(x) D. r(x) / e(x) = p(x)

OpenStudy (alexandervonhumboldt2):

p(x)=r(x)-e(x) i think

OpenStudy (anonymous):

r(x)-e(x)=p(x)?

OpenStudy (alexandervonhumboldt2):

i think yes but not sure

OpenStudy (anonymous):

oh okay. so should i put that down